# Efficient Visual Transformer by Learnable Token Merging

### v1

@@ 4.2 Object Detection
Results. We adopt a comparative study of our LTM Transformers against other lightweight feature backbones within the SSDLite object detection framework. The results, as detailed in Table III, illustrate significant improvements in object detection performance when the feature backbone is upgraded to include LTM-Transformer blocks. For example, substituting MobileViT-S with LTM-MobileViT-S enhances the mAP by $0.7\%$ while concurrently reducing FLOPs by $0.3$G. Additionally, SSDLite equipped with LTM-EfficientViT achieves a substantial performance increase of $6.9\%$ while maintaining the same FLOPs as MobileNetV3.
4.3 Instance Segmentation
In this section, we assess the efficacy of LTM when applied to instance segmentation tasks using the COCO dataset [42]. We utilize Mask R-CNN [49] equipped with a Feature Pyramid Network (FPN) as the segmentation head, built on the LTM-EfficientViT-B1 feature backbone. For comparative analysis, we include EfficientViT-B1 [7] and EViT [21] as baseline models. Both our models and the baselines are trained on the training split of the COCO dataset and evaluated on the validation split, adhering to the protocols established by [50]. The training duration is set to 12 epochs, consistent with the 1$\times$ schedule described in [50]. The AdamW optimizer is employed for training following the practices of [21]. We initiate the learning rate at $0.001$, which is then gradually reduced following a cosine learning rate schedule. Performance metrics reported include the mean bounding box Average Precision (mAPb) and mean mask Average Precision (mAPm), along with bounding box Average Precision (APb) and mask Average Precision (APm) at IoU thresholds of 0.5 and 0.75. The findings, detailed in Table IV, demonstrate that LTM-EfficientViT-B1 consistently enhances segmentation performance across various thresholds.
@@ 4.4 Ablation Study
Study on the effects of LTM in reducing IB loss. We study the effectiveness of LTM-Transformer in reducing the IB loss and the variational upper bound of IB loss, which is the IB bound, across three visual transformers, including MobileViT-S, MobileViT-XS, and EfficientViT (r224). We compare the performance of the visual transformers with the baseline token merging method, ToME [14], and the corresponding LTM-Tranformer models with all the tansformer blocks replaced with the LTM-Transformer blocks. The results are shown in Table V. The results indicate that although ToMe reduces the IB loss and the IB bound, thereby adhering to the IB principle which aims to enhance the correlation of features with class labels while reducing their correlation with the input, LTM can further decrease the IB loss and IB bound. In particular, our LTM-Transformer models improve the vanilla visual transformers and the ToMe models by a large margin in terms of both IB loss and top-1 accuracy.
Model # Params FLOPs Top-1 IB Bound IB Loss
MobileViT-S 5.6 M 1.40 G 78.4 0.05782 -0.00432
ToMe-MobileViT-S 5.6 M 1.22 G 76.7 0.04542 -0.00913
LTM-MobileViT-S 5.9 M 1.17 G 79.7 0.02791 -0.01773
MobileViT-XS 2.3 M 0.71 G 74.8 0.05539 -0.00419
ToMe-MobileViT-XS 2.3 M 0.54 G 72.7 0.04583 -0.00647
LTM-MobileViT-XS 2.5 M 0.52 G 75.8 0.03082 -0.01618
EfficientViT (r224) 9.1 M 0.52 G 79.4 0.06014 -0.00451
ToMe-EfficientViT (r224) 9.1 M 0.47 G 78.8 0.04642 -0.00732
LTM-EfficientViT (r224) 9.5 M 0.44 G 80.2 0.02886 -0.01539
@@
TABLE V: Ablation Study on the effects of LTM in reducing IB loss.
@@
Methods # Params FLOPs Training Time (Hours) Top-1
MobileViT-XS 2.3 M 0.70 G 73.5 75.8
ToMe-MobileViT-XS 2.3 M 0.54 G 73.5 72.7
ToFu-MobileViT-XS 2.3 M 0.54 G 73.5 73.3
LTMP-MobileViT-XS 2.3 M 0.56 G 73.8 73.9
LTM-MobileViT-XS 2.5 M 0.52 G 91.0 76.8
MobileViT-S 5.6 M 1.40 G 89.5 78.4
ToMe-MobileViT-S 5.6 M 1.22 G 89.5 76.7
ToFu-MobileViT-S 5.6 M 1.22 G 89.5 77.2
LTMP-MobileViT-S 5.6 M 1.17 G 90.0 77.5
LTM-MobileViT-S 5.9 M 1.22 G 105.0 79.7
EfficientViT-B1 [r224] 9.1 M 0.52 G 73.0 79.4
ToMe-EfficientViT-B1 [r224] 9.1 M 0.47 G 73.0 78.8
ToFu-EfficientViT-B1 [r224] 9.1 M 0.47 G 73.0 79.0
LTMP-EfficientViT-B1 [r224] 9.1 M 0.50 G 73.3 79.2
LTM-EfficientViT-B1 [r224] 9.5 M 0.44 G 91.0 80.2
EfficientViT-B1 [r288] 9.1 M 0.86 G 95.5 80.4
ToMe-EfficientViT-B1 [r288] 9.1 M 0.73 G 95.5 79.7
ToFu-EfficientViT-B1 [r288] 9.1 M 0.73 G 95.5 89.8
LTMP-EfficientViT-B1 [r288] 9.1 M 0.76 G 95.9 80.0
LTM-EfficientViT-B1 [r288] 9.5 M 0.70 G 110.5 81.0
TABLE VI: Training time (minutes/epoch) comparisons between LTM-Transformers and their baseline models.
Study on the IB loss and IB bound at different layers. To study how the IB loss $\textup{IB}(G)$, and the variational upper bound for the IB loss, $\textup{IBB}(G)$, decrease with respect to layer index $\ell$ of a LTM-Transformer network, $\textup{IB}(G)$ and $\textup{IBB}(G)$ for both MobileViT-S and LTM-MobileViT-S across different transformer layers are illustrated in Figure 3. Both models contain 9 transformer layers. It is observed from Figure 3 that both $\textup{IB}(G)$ and $\textup{IBB}(G)$ decrease in deeper layers with larger layer indices of MobileViT-S and LTM-MobileViT-S. This observation suggests that features in deeper layers correlate more closely with the class labels and less with the input features, adhering to the IB principle. Moreover, LTM-MobileViT-S reduces both $\textup{IB}(G)$ and $\textup{IBB}(G)$ to lower levels in deeper layers compared to MobileViT-S. These observations evidence that the mask module in the LTM-Transformer block which generates the informative token merging task by (3.2) can effectively reduce both $\textup{IB}(G)$ and $\textup{IBB}(G)$, better adhering to the IB principle than the vanilla MobileViT-S. At the early stage of the training after $100$ epochs, the IB bound and the IB loss of LTM-MobileViT-S are similar to those of the MobileViT-S. After training for $300$ epochs, the IB bound and the IB loss of LTM-MobileViT-S are much smaller than those of the MobileViT-S.
@@
5 More Ablation Studies
We conduct more ablation studies in this section, including the training time comparison, the training loss and test loss of LTM-Transformers, and the visualization results illustrating the effectiveness of a LTM-Transformer in selecting informative tokens during the token merging process. We compare the training time of LTM-Transformer models with the competing baselines for token merging in Table VI in Section 5.1. Figure 4 in Section 5.2 illustrates the training loss and the test loss during the training process of LTM-Transformers and their baselines, highlighting that the test loss of the LTM-Transformer networks exhibits a more rapid decline compared to that of their baselines.
5.1 Training Time Evaluation
We evaluate the training cost of our LTM-Transformer models and the baseline models on the training set of ImageNet-1k. The training is performed on 4 NVIDIA A100 GPUs with an effective batch size of 512 images. We report the overall training time of 300 epochs. We also include the training time of ToMe [14], ToFu [12], and LTMP [13] for comparison. It is noted that ToMe, ToFu, and LTMP are applied to pre-trained models. Therefore, the training time for ToMe, ToFu, and LTMP includes the training time of the baseline models. In contrast, our models are trained from scratch. The training time of various models are shown in Table VI. The training overhead of LTM-Transformers mainly comes from the computation of $\left\{\phi(\tilde{X}_{i},a)\right\}_{i\in{\cal B}_{j},a\in[C]}$, $\left\{Q^{(t-1)}(\tilde{X}\in a|Y=y)\right\}_{a\in[C],y\in[C]}$, and $\left\{\tilde{\cal C}^{(t-1)}_{a}\right\}_{a=1}^{C}$ as described in Algorithm 1. It is observed from the Table VI that the training time of LTM models is comparable to the training time of the competing token merging methods. In addition, LTM largely resolves the issue of significant prediction accuracy drops after token merging by ToMe, ToFu, and LTMP.
5.2 Training Loss and Test Loss of LTM-Transfomers
@@
In this section, we illustrate the training loss and the test loss of LTM-MobileViT-XS, LTM-MobileViT-S, and LTM-EfficientViT-B1. In comparison, we also illustrate the training loss and test loss of MobileViT-XS, MobileViT-S, and EfficientViT-B1. All the models are trained for 300 epochs. The plots are illustrated in Figure 4. It can be observed that LTM-Transformer networks achieve lower training losses and test losses at the end of the training, which demonstrates the benefit of LTM-Transformers in improving the performance of the visual transformers through the IB-inspired token merging.
Figure 5: Visualization of merging weights in the first LTM-Transformer block in LTM-MobileViT-S.
5.3 Visualization Results

### v2

@@ 4.1 Image Classification
4.2 Object Detection
@@
Results. We conduct a comparative study of our LTM Transformers against other lightweight feature backbones within the SSDLite object detection framework. The results, as detailed in Table III, illustrate significant improvements in object detection performance when the feature backbone is upgraded to include LTM-Transformer blocks. For example, substituting MobileViT-S with LTM-MobileViT-S enhances the mAP by $0.7\%$ while concurrently reducing FLOPs by $0.3$G. In addition, SSDLite equipped with LTM-EfficientViT achieves a substantial performance increase of $6.9\%$ while maintaining the same FLOPs as MobileNetV3.
4.3 Instance Segmentation
In this section, we assess the efficacy of LTM when applied to instance segmentation tasks using the COCO dataset [45]. We utilize Mask R-CNN [46] equipped with a Feature Pyramid Network (FPN) as the segmentation head, built on the LTM-EfficientViT-B1 feature backbone. For comparative analysis, we include EfficientViT-B1 [6] and EViT [21] as baseline models. Both our models and the baselines are trained on the training split of the COCO dataset and evaluated on the validation split, adhering to the protocols established by [47]. The training duration is set to 12 epochs, consistent with the 1$\times$ schedule described in [47]. The AdamW optimizer is employed for training following the practices of [21]. We initiate the learning rate at $0.001$, which is then gradually reduced following a cosine learning rate schedule. Performance metrics reported include the mean bounding box Average Precision (mAPb) and mean mask Average Precision (mAPm), along with bounding box Average Precision (APb) and mask Average Precision (APm) at IoU thresholds of 0.5 and 0.75. The findings, detailed in Table IV, demonstrate that LTM-EfficientViT-B1 consistently enhances segmentation performance across various thresholds. For example, LTM-EfficientViT-B1 outperforms EfficientViT-B1 by $0.8\%$ and $0.9\%$ in mAPb and mAPm, respectively.
@@ 4.4 Ablation Study
4.4.1 Study on the Impact of Compression Ratio
We conduct an ablation study on the compression ratio of token merging on ViT-B. We train LTM-ViT-B models with different compression ratios in both the train-from-scratch setup and fine-tuning setup. In the fine-tuning setup, all models are fine-tuned for 50 epochs. It is observed from the results in Table V that although a smaller compression ratio can result in a slight accuracy drop, the LTM-ViT-B with a compression ratio of 0.65 can still achieve better performance than the original ViT-B model either fine-tuned from the pre-trained checkpoint or trained from scratch.
Methods FLOPs (G) $r$ Train-from-scratch Fine-tuning
ViT-B 17.58 1.00 83.74 83.74
LTM-ViT-B 16.55 0.95 84.43 84.32
LTM-ViT-B 15.25 0.90 84.46 84.29
LTM-ViT-B 14.19 0.85 84.33 84.35
LTM-ViT-B 14.89 0.80 84.15 84.20
LTM-ViT-B 13.49 0.75 83.95 84.05
LTM-ViT-B 12.85 0.70 83.87 83.96
LTM-ViT-B 11.95 0.65 83.75 83.81
LTM-ViT-B 11.03 0.60 83.53 83.56
LTM-ViT-B 10.15 0.55 83.07 83.14
LTM-ViT-B 9.63 0.50 82.87 82.95
LTM-ViT-B 8.77 0.45 83.46 83.51
LTM-ViT-B 8.30 0.40 82.23 82.37
TABLE V: Performance comparison between LTM-ViT-B with different compression ratios. The models in the fine-tuning setup are fine-tuned for 50 epochs.
4.4.2 Study on the Effects of LTM in Reducing IB Loss
@@
We study the effectiveness of LTM in reducing the IB loss and the variational upper bound of IB loss, which is the IB bound, across three vision transformers, including MobileViT-S, MobileViT-XS, and EfficientViT (r224). We compare the performance of the vision transformers with the baseline token merging method, ToME [15], LTMP [14], and the corresponding LTM-Transformer models with all the transformer blocks replaced with the LTM blocks. The ablation study results for the fine-tuning setup are shown in Table VI. The ablation study results for the train-from-scratch setup are shown in Table VII. The results indicate that although ToMe and LTMP reduce the IB loss and the IB bound in both the fine-tuning setup and the train-from-scratch setup, thereby adhering to the IB principle, which aims to enhance the correlation of features with class labels while reducing their correlation with the input, LTM can further decrease the IB loss and IB bound. In particular, our LTM models improve the vanilla vision transformers, the ToMe models, and the LTMP models by a large margin in terms of both IB loss and top-1 accuracy for both the fine-tuning setup and the train-from-scratch setup. For instance, the LTMP-ViT-B fine-tuned for 50 epochs reduces the IB loss of ViT-B by $0.00333$. The LTM-ViT-B fine-tuned for 50 epochs further reduces the IB loss of LTMP-ViT-B fine-tuned for 50 epochs by $0.00866$.
Model FLOPs Top-1 IB Bound IB Loss
0 1 10 50 0 1 10 50 0 1 10 50
MobileViT-S 1.40 G 78.40 - - - 0.05782 - - - -0.00432 - - -
ToMe-MobileViT-S 1.22 G 76.72 - - - 0.04931 - - - -0.00525 - - -
LTMP-MobileViT-S 1.26 G - 77.53 77.82 78.14 - 0.04902 0.04735 0.04542 - -0.00765 -0.00874 -0.00913
LTM-MobileViT-S 1.17 G - 77.72 78.34 79.05 - 0.03095 0.02967 0.02683 - -0.01430 -0.01576 -0.01692
EfficientViT-B1 0.52 G 79.40 - - - 0.06014 - - - -0.00451 - - -
ToMe-EfficientViT-B1 0.47 G 78.81 - - - 0.04642 - - - -0.00732 - - -
LTMP-EfficientViT-B1 0.52 G - 79.21 79.32 79.40 - 0.04537 0.04219 0.03970 - -0.00802 -0.00916 -0.00995
LTM-EfficientViT-B1 0.44 G - 79.39 79.62 80.22 - 0.02874 0.02703 0.02635 - -0.01585 -0.01664 -0.01704
ViT-B 17.58 G 83.74 - - - 0.05539 - - - -0.00419 - - -
ToMe-ViT-B 13.12 G 82.86 - - - 0.04583 - - - -0.00647 - - -
LTMP-ViT-B 13.46 G - 83.29 83.44 83.55 - 0.04392 0.04275 0.04086 - -0.00665 -0.00693 -0.00752
LTM-ViT-B 12.85 G - 83.35 83.76 83.96 - 0.03732 0.03506 0.03082 - -0.01425 -0.01572 -0.01618
@@
TABLE VI: Ablation study on the effects of LTM in reducing IB loss in the fine-tuning setup. Both LTMP models and LTM models fine-tuned for 1, 10, 50 epochs are evaluated.
Model FLOPs Top-1 IB Bound IB Loss
MobileViT-S 1.40 G 78.40 0.05782 -0.00432
ToMe-MobileViT-S 1.22 G 76.72 0.04931 -0.00525
LTMP-MobileViT-S 1.26 G 78.14 0.04542 -0.00913
LTM-MobileViT-S 1.17 G 79.68 0.02425 -0.01725
EfficientViT-B1 0.52 G 79.40 0.06014 -0.00451
ToMe-EfficientViT-B1 0.47 G 78.81 0.04642 -0.00732
LTMP-EfficientViT-B1 0.52 G 79.40 0.03970 -0.00995
LTM-EfficientViT-B1 0.44 G 80.20 0.02689 -0.01730
ViT-B 17.58 G 83.74 0.05539 -0.00419
ToMe-ViT-B 13.12 G 82.86 0.04583 -0.00647
LTMP-ViT-B 13.46 G 83.55 0.04086 -0.00752
LTM-ViT-B 12.85 G 83.87 0.03094 -0.01636
TABLE VII: Ablation study on the effects of LTM in reducing IB loss in the train-from-scratch setup.
@@
4.4.3 Study on the IB loss and IB bound at Different Layers
To study how the IB loss $\textup{IB}(G)$, and the variational upper bound for the IB loss, $\textup{IBB}(G)$, decrease with respect to layer index $\ell$ of a LTM-Transformer network, $\textup{IB}(G)$ and $\textup{IBB}(G)$ for both MobileViT-S and LTM-MobileViT-S across different transformer layers are illustrated in Figure 2. Both models contain 9 transformer layers. It is observed from Figure 2 that both $\textup{IB}(G)$ and $\textup{IBB}(G)$ decrease in deeper layers with larger layer indices of MobileViT-S and LTM-MobileViT-S. This observation suggests that features in deeper layers correlate more closely with the class labels and less with the input features, adhering to the IB principle. Moreover, LTM-MobileViT-S reduces both $\textup{IB}(G)$ and $\textup{IBB}(G)$ to lower levels in deeper layers compared to MobileViT-S. These observations evidence that the mask module in the LTM-Transformer block, which generates the informative token merging mask by (3) can effectively reduce both $\textup{IB}(G)$ and $\textup{IBB}(G)$, better adhering to the IB principle than the vanilla MobileViT-S. At the early stage of the training after $100$ epochs, the IB bound and the IB loss of LTM-MobileViT-S are similar to those of the MobileViT-S. After training for $300$ epochs, the IB bound and the IB loss of LTM-MobileViT-S are much smaller than those of the MobileViT-S.
Figure 4: Comparisons of token merging weights generated in the first transformer block in LTM-MobileViT-S and LTMP-MobileViT-S. The target tokens resulting from the token merging process are marked with red boxes in the first row. The token merging weights for tokens merged into the target tokens by LTMP-MobileViT-S are illustrated in the second row. The token merging weights for tokens merged into the target tokens by LTM-MobileViT-S are illustrated in the third row. Tokens that are more semantically similar to the target tokens should receive higher merging weights so as to achieve informative token merging. Examples of the undesirable tokens, which are not semantically similar to the target tokens, are marked with white diamonds. In contrast with LTMP, the heatmaps of the merging weights generated by LTM illustrate that LTM usually assigns larger weights to tokens that are more semantically similar to the target tokens.
4.4.4 Training Loss and Test Loss of LTM-Transformers
@@
In this section, we illustrate the training loss and the test loss of LTM-MobileViT-XS, LTM-MobileViT-S, and LTM-EfficientViT-B1. In comparison, we also illustrate the training loss and test loss of MobileViT-XS, MobileViT-S, and EfficientViT-B1. All the models are trained for 300 epochs for each cycle with two cycles. The plots are illustrated in Figure 3 for the second cycle. It can be observed that LTM-Transformer networks achieve lower training losses and test losses at the end of the training, which demonstrates the benefit of LTM-Transformers in improving the performance of the visual transformers through the IB-inspired token merging.
4.4.5 Visualization Results